OpenStudy (anonymous):

prove sin^2(x)+cos^2(x)=1

OpenStudy (anonymous):

this is for mimi too hard for me

OpenStudy (mimi_x3):

lols , im not that stupid againss to prove it

OpenStudy (anonymous):

NO YOU PROVE it

OpenStudy (anonymous):

prove it what the....

OpenStudy (anonymous):

please mimi help prove this identity

OpenStudy (mimi_x3):

LOLS

OpenStudy (anonymous):

mimi i m serious :(

OpenStudy (mimi_x3):

yeah right , im not gullible

OpenStudy (anonymous):

plz help and ishaan u and hahd too

OpenStudy (anonymous):

she's not falling for it

OpenStudy (anonymous):

lana is here too help

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!! help

OpenStudy (lalaly):

sin t = y/r, cos t = x/r, and r^2 = x^2+y^2. so sin^2x = y^2/r^2 and cos^2x = x^2/r^2. (y^2/r^2) + (x^2/r^2). (y^2+x^2)/r^2 = r^2/r^2 = 1.

OpenStudy (lalaly):

lolz

OpenStudy (anonymous):

explain plz

OpenStudy (anonymous):

i cant understand it !!!!

OpenStudy (lalaly):

i cant im busy !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

mimi will

OpenStudy (anonymous):

medal is guranteed

OpenStudy (anonymous):

is this a serious question? what is your definition of sine and cosine? the usual one is points on the unit circle in which case there is nothing to prove

OpenStudy (mimi_x3):

(b/c)^2 + (a/c)^2 a^2/c^2 + a^2/c^2 = b^2 +a^2/c^2 c^2/c^2 = 1

OpenStudy (anonymous):

mimi and angela help

OpenStudy (mimi_x3):

lols , im jst bored thats why im answering this stupid question lmao

OpenStudy (anonymous):

it is not stupid

OpenStudy (anonymous):

mimi u r stupid

OpenStudy (mimi_x3):

yes it is (Y))

OpenStudy (mimi_x3):

excuse me , how rude :P

OpenStudy (anonymous):

triangle with sides a,b Take sin and cos, square, add = 1

OpenStudy (anonymous):

and u r x3 times stupid

OpenStudy (mimi_x3):

excuse me , yur such a noob x3 is not times 3 :P

OpenStudy (anonymous):

is is

OpenStudy (mimi_x3):

huh ?

OpenStudy (anonymous):

x3=3 times mimix3 *3=mimi x9

OpenStudy (mimi_x3):

lolss , such a noob x3 is not times 3 arghh

OpenStudy (anonymous):

hehehehe jst jokin chill out

OpenStudy (mimi_x3):

i know manss :P

OpenStudy (mimi_x3):

im not stupid , im so~ bored *sigh*

OpenStudy (mimi_x3):

lmao :L:L:L